OpenStudy (anonymous):

The distance between stars is measured in light-years—the distance that light travels in a year. One of the closest stars to our Sun is Alpha Centauri; another is Sirius. Sirius is 4.21 light-years farther away from our Sun than Alpha Centauri. Sirius is also twice as far as the distance between Alpha Centauri and the Sun, less 0.08 light years. This situation is expressed below as an equation, where x is the distance in light-years from our Sun to Alpha Centauri. x + 4.21 = 2(x − 0.08) How many light-years is it from our Sun to Alpha Centauri? A. 1.43 B. 4.13 C.

OpenStudy (anonymous):

c.4.29 d.4.37

OpenStudy (anonymous):

x + 4.21 = 2(x − 0.08) U need to solve for x

OpenStudy (anonymous):

ummmm how

OpenStudy (anonymous):

x + 4.21 = 2(x − 0.08) x+4.21 =2x - 0.16 x+4.21 +0.16 = 2x-0.16 +0.16 x+4.37 = 2x x+4.37 -x =2x-x x=4.37

OpenStudy (anonymous):

Did u understand what I did above ?

OpenStudy (anonymous):

omg i get it now we did this in place thnxs

OpenStudy (anonymous):

welcome